OpenStudy (anonymous):

Find the first, second, and third derivatives of the function: f(x) = x^3. Please show all work, I've been having trouble with following the steps that require me to divide any variables that have "delta x" or "delta^2" joined with them.

OpenStudy (anonymous):

f(x)=x3 f'(x)=3x2 f''(x)=6x f'''(x)=6 you use the formula for differentiation d/dx(x^n)=nx^(n-1)

OpenStudy (anonymous):

Thanks Sarkar, does this formula work when trying to find the derivative of all nonlinear functions? Or is this just a special case? Say, if I were to try and find the derivatives of f(x) = x^2 would the same formula apply?

OpenStudy (anonymous):

yes,this is a general formula...it WILL work in all cases having such a configuration x^n

OpenStudy (anonymous):

I see, thanks for the help; it's very much appreciated